# XENONnT WIMP Search: Signal & Background Modeling and Statistical Inference

### v1

@@ Abstract
The XENONnT experiment searches for weakly-interacting massive particle (WIMP) dark matter scattering off a xenon nucleus. In particular, XENONnT uses a dual-phase time projection chamber with a 5.9-tonne liquid xenon target, detecting both scintillation and ionization signals to reconstruct the energy, position, and type of recoil. A blind search for nuclear recoil WIMPs with an exposure of 1.1 tonne-years yielded no signal excess over background expectations, from which competitive exclusion limits were derived on WIMP-nucleon elastic scatter cross sections, for WIMP masses ranging from 6 GeV/ up to the TeV/ scale. This work details the modeling and statistical methods employed in this search. By means of calibration data, we model the detector response, which is then used to derive background and signal models. The construction and validation of these models is discussed, alongside additional purely data-driven backgrounds. We also describe the statistical inference framework, including the definition of the likelihood function and the construction of confidence intervals.
@@ I Introduction
Astrophysical and cosmological observations at various scales provide clear evidence for the existence of dark matter (DM) as a fundamental building block of our universe [1, 2]. Numerous extensions of the standard model of particle physics predict additional fundamental particles as potential candidates for DM [3, 4]. Measuring the direct interaction of particles from the DM halo of the Milky Way with a suitable detector could provide conclusive evidence for the particle DM hypothesis. The XENONnT experiment aims to detect a signal from weakly interacting massive particles (WIMPs) elastically scattering off xenon nuclei with a detector operated underground at the INFN Laboratori Nazionali del Gran Sasso (LNGS) in Italy. The collaboration published the first WIMP search results from a data-taking period between July and November 2021 with a total live time of days and a fiducial mass of , referred to as science run 0 (SR0). We performed a blind analysis and observed no significant excess above background expectations [5]. This led to a minimum upper limit on the spin-independent WIMP-nucleon cross section of for a WIMP mass of at confidence level.
@@
The detector used for the XENONnT experiment is a dual-phase xenon time projection chamber (TPC). A particle interaction in the active liquid xenon (LXe) target results in prompt vacuum ultraviolet scintillation light as well as free ionization electrons. The instantaneous light signal, called the S1 signal, is detected by two arrays of photomultiplier tubes (PMTs) at the top and bottom of the approximately cylindrical TPC. An electric field ( average electric drift field during SR0) is applied to the target volume to drift the ionization electrons toward the liquid xenon surface. Here, they are extracted and accelerated into a xenon gas volume above the liquid via a stronger electric field ( electric extraction field during SR0). The kinetic energy of the accelerated electrons in the gas phase is sufficient for the emission of electroluminescent light, which is proportional to the number of extracted electrons. This second light signal is referred to as the S2 signal. The distribution of the S2 signal detected by the PMTs in the top array is used to infer the position of the initial interaction in the horizontal plane parallel to the liquid surface, which defines the radial coordinate . The vertical component follows from the drift time of electrons, which is determined as the time difference between the S1 and S2 signals. This three-dimensional position reconstruction of the interaction vertex enables the discrimination of multi-site events as well as the selection of an inner fiducial volume (FV) within the TPC. This volume benefits from a particularly low background level thanks to the excellent self-shielding properties of LXe, given its high density. Moreover, the relative magnitudes of the S1 and S2 signals can be used to determine if a particle interacted with the xenon nucleus (expected from WIMPs) or xenon shell electrons (typical of the dominant sources of background), i.e. nuclear recoil (NR) or electronic recoil (ER) events. The electric drift and extraction fields are established by three parallel-wire electrodes (cathode, gate, and anode from bottom to top). To reduce wire sagging, two and four horizontal perpendicular wires support the gate and anode electrodes, respectively.
@@
The TPC is nested in an active neutron veto, which in turn is housed in an active muon veto. Both veto systems are water Cherenkov detectors. More details on the TPC, the veto systems, and other subsystems of the detector are provided in [6, 7].
@@ II Modeling the detector response
Modeling the detector response to energy depositions allows for a powerful discrimination between NR signal and ER background in S1–S2 space. NR events are produced by particles scattering elastically off xenon nuclei, while in ER events, particles scatter off xenon shell electrons. The different energy loss processes involved after the two types of recoil cause ER and NR events to form separate populations in S1–S2 space, which is exploited in the WIMP DM search.
In this section, we describe the TPC response model to energy depositions in LXe via ER or NR interactions, which is obtained from fits to calibration data. This detector response model is separated into two parts. The first one is the empirically parametrized LXe emission model, which describes the production processes of the detectable quanta, i.e. the conversion of deposited energy into scintillation photons and ionization electrons. The second part is the detector reconstruction model, which covers the conversion from the produced photons and electrons into the observed and spacetime-dependence corrected S1 and S2 signals (cS1 and cS2). Due to its complexity, it is unfeasible in the fit to simulate processes of photon and electron propagation on a per-quantum basis. Instead, toy-Monte-Carlo (toy-MC) simulations of the detector reconstruction model are used which sample from effective maps provided either by data-driven analyses described in [10] or by simulation-driven analyses using the XENONnT Monte Carlo (MC) framework [11, 12]. Then all the model parameters in the simulations are fit to both ER and NR calibration data. The two parts of the detector response model together with the toy-MC simulation workflow are described in Secs. II.1 and II.2, and the fit to ER and NR calibration data is detailed in Sec. II.3.
@@ II.1 Liquid Xenon Emission Model
The production of quanta from energy depositions in a xenon target is complex and lacks a comprehensive description derived from first principles. Thus, a semi-empirical parametrization of the emission model is commonly used. The parametrization used in XENONnT SR0 is similar to XENON1T [9], which is based on the Noble Element Simulation Technique (NEST) model described in [13, 14]. The simulation workflow of the emission model is described in detail in Appendices A and B, and is summarized in the following.
@@
In the region of interest (ROI) for the SR0 WIMP search, in [0, 100] PE and in [, ] PE, NR interactions have a larger recombination probability than ER interactions given the same total energy deposition. This leads to smaller ratios of S2 to S1 areas, which is the principle behind ER-NR discrimination in LXe TPCs. We parametrize the mean recombination fraction for ER interactions using the Thomas-Imel box model [15] with an additional Fermi-Dirac term, as in [9]. For NR interactions, the emission model follows the parametrization used in the NEST v1 model detailed in [14]. While there are newer versions of fits available in the NEST framework that use a different parametrization for NR interactions (which we refer to as NEST v2 in the following), we choose to stay with the previous version (referred to as NEST v1). The model is simpler, fits our data well, and the best-fit values and uncertainties of parameters of NEST v2 were not published at the time of the analysis.
@@ II.2 Detector Reconstruction Model
For S1 signals, the spatial dependence of geometric effects during photon propagation, the photon detection efficiency of the PMTs, and the effect of double photoelectron emission (DPE) from the PMT photocathode [16, 17] are all modeled in the simulations. For S2 signals, we model the loss of electrons along their drift path due to attachment to electronegative impurities as well as due to electric field effects close to the TPC wall, the efficiency of electron extraction from the liquid into the gas phase, and the single-electron gain of extracted electrons to detected photoelectrons (PE) via the electroluminescence process in gas. In addition to these physical processes, we also account for the influence of software reconstruction effects on the signals. For both signal types, the software reconstruction process can introduce biases and fluctuations in the recorded signal size. We also account for the S1 signal reconstruction efficiency, which vanishes at about 3 PE, as we require that at least 3 PMTs detect a photon within 50 ns around the maximal amplitude. Finally, we model the uncertainty of the event position reconstruction and the acceptances of data quality selections [10].
@@
The NR model is calibrated using an external 241AmBe neutron source (referred to as AmBe source in the following). It emits a fast neutron via the 9BeC capture reaction. Neutrons often scatter multiple times in the LXe target in so-called multi-scatter (MS) events, creating distinct energy depositions at each interaction site, thus the topologies of neutron events must be considered in the modeling. Since the neutrons travel at about 10 % of the speed of light and the S1 width is , the S1 signals from the separate energy depositions in a MS event get reconstructed into one merged S1 signal. The S2 signals might be distinguished based on the separation in spacetime between the energy depositions of the individual scatters, but can get (partially) merged if the interaction sites are close together. Due to the low drift field of in XENONnT SR0, the separation resolution between two S2 signals is reduced with respect to the design because of the increased drift time and consequently larger S2 spread due to the diffusion of electrons. This effect is accounted for in the fit of the NR model to the neutron calibration data: we first produce the photons and electrons for each energy deposition in a MS event separately, based on the emission model. While the primary scintillation photons are summed to produce the merged S1 signal, we only sum S2 signals that will become part of the largest S2 signal after going through the remaining detector reconstruction processes. This information is provided by PMT waveform simulations of neutron scatters from the AmBe source, using the Monte Carlo (MC) framework of XENONnT [11, 12] (see also Sec. III.3 for details on neutron simulations). The results of these waveform simulations are used as inputs to the toy-MC simulations and fitting framework. Data selections that remove MS events (both resolved and unresolved) are directly applied to these inputs, because their acceptances can only be determined from full simulations, and are correspondingly dropped from the selection acceptance curves for the NR fit.
@@ II.3 Fit to Calibration Data
The parameters in the full detector response model are fitted to 220Rn, 37Ar, and AmBe calibration data. The ROI selection is the same as the SR0 WIMP search region defined in Sec. II.1. All data quality selections are applied to the calibration data within the ROI. For the AmBe neutron data, an additional selection is applied using coincident gammas detected in the neutron veto. In 9BeC capture reactions in the AmBe source, 12C is left in the first excited state after neutron emission in about 50 % of the cases, estimated with neutron veto data [20]. The 12C excited state promptly decays via the emission of a 4.44 MeV gamma ray, coincident with the neutron emission. The requirement of the time coincidence within between the 4.44 MeV gamma ray detected in the neutron veto and the event in the TPC, leads to a clean calibration dataset, with only about one expected event from accidental coincidences between the two detectors in the dataset (0.05 % of the resulting sample). The events recorded in the ROI that pass (fail) the neutron veto coincidence selection are shown in red (gray) in Fig. 1.
@@
In the ER emission model, the parametrization is the same as in XENON1T [9] with very wide flat priors, referred to as free priors. For the NR emission model, the parameter priors are taken from [14]. A Gaussian-like distribution with different widths on either side is used as the prior of parameters with asymmetric uncertainties. However, due to the low drift field of 23 V/cm, where no literature data on NR yields exist, the validity of the field dependence in the model is unverified. Therefore, the field dependence in the emission model from [14] was modified. Two scaling parameters were freed ( and in Eq. 24 and Eq. 27 in the appendix), and two field dependence parameters ( and ) were fixed to the reported best-fit values. The widths of all other parameter priors were doubled in order to allow more freedom in the fit. Because the dependence of NR signals on the drift field has been shown to be small (see also literature measurements shown in Fig. 3), we consider this a reasonable choice.
The fit results of all emission model parameters are summarized in Tab. 2 in the appendix. No significant tension between the posteriors and the priors of any parameter is found. Suitable goodness-of-fit (GOF) tests are performed to assess whether the best-fit models adequately describe the 220Rn and AmBe calibration data. Specifically, we employ binned Poisson likelihood tests in the – space. We adopt an equiprobable binning scheme using the GOFevaluation package [37], ensuring that the number of expected events under the best-fit model is the same in each bin, . In Fig. 2 the results of the 2D – Poisson likelihood test are shown for both the ER and the NR fit, overlaid with the 220Rn and the AmBe calibration data, respectively. The resulting p-values of the tests indicate no significant discrepancy between the best-fit models and the calibration data.
@@ III.1 WIMP Signal Model
The axial-vector structure factor of xenon is taken from [47]. In the XENONnT SR0 analysis, the mean of the structure factor is used and the uncertainty is neglected since it is much smaller than the uncertainty from the NR fit.
@@ III.2 Electronic Recoil Background Model
ER events are one of the dominant background sources in the XENONnT SR0 WIMP search, primarily originating from -decays of intrinsic radioactive contaminants such as 214Pb (a product of the 222Rn decay chain) and 85Kr. Contributions from 136Xe double- decays, solar neutrino interactions, and gamma events from the materials are also taken into account. In the ROI for the WIMP search, the total ER energy spectrum is approximately flat. The distribution of ER background events in – space is generated from the ER model fitted to 220Rn and 37Ar calibration data, as discussed in Sec. II.3. The total ER rate is left unconstrained and is fitted in the final WIMP likelihood, which will be discussed in Sec. IV.
@@
along each principal component is used to quantify its importance. Here, and are the probability of a 50 GeV WIMP and ER background in the bin of cS1–cS2 space, respectively. A larger variance of means that the uncertainty of that component can be more impactful for the WIMP sensitivity. In the end, the two most important principal components are kept as nuisance parameters. Their shape uncertainty, constrained by the calibration data, is shown in Fig. 6.

### v2

@@ Abstract
The XENONnT experiment searches for weakly-interacting massive particle (WIMP) dark matter scattering off a xenon nucleus. In particular, XENONnT uses a dual-phase time projection chamber with a 5.9-tonne liquid xenon target, detecting both scintillation and ionization signals to reconstruct the energy, position, and type of recoil. A blind search for nuclear recoil WIMPs with an exposure of 1.1 tonne-years ( fiducial mass) yielded no signal excess over background expectations, from which competitive exclusion limits were derived on WIMP-nucleon elastic scatter cross sections, for WIMP masses ranging from 6 GeV/ up to the TeV/ scale. This work details the modeling and statistical methods employed in this search. By means of calibration data, we model the detector response, which is then used to derive background and signal models. The construction and validation of these models is discussed, alongside additional purely data-driven backgrounds. We also describe the statistical inference framework, including the definition of the likelihood function and the construction of confidence intervals.
@@ I Introduction
Astrophysical and cosmological observations at various scales provide clear evidence for the existence of dark matter (DM) as a fundamental building block of our universe [1, 2]. Numerous extensions of the standard model of particle physics predict additional fundamental particles as potential candidates for DM [3, 4]. Measuring the direct interaction of particles from the DM halo of the Milky Way with a suitable detector could provide conclusive evidence for the particle DM hypothesis. The XENONnT experiment aims to detect a signal from weakly interacting massive particles (WIMPs) elastically scattering off xenon nuclei with a detector operated underground at the Istituto Nazionale di Fisica Nucleare (INFN) Laboratori Nazionali del Gran Sasso (LNGS) in Italy. The collaboration published the first WIMP search results from a data-taking period between July and November 2021 with a total live time of days and a fiducial mass of , referred to as science run 0 (SR0). We performed a blind analysis and observed no significant excess above background expectations [5]. This led to a minimum upper limit on the spin-independent WIMP-nucleon cross section of for a WIMP mass of at confidence level (CL).
@@
The detector used for the XENONnT experiment is a dual-phase xenon time projection chamber (TPC). A particle interaction in the active liquid xenon (LXe) target results in prompt vacuum ultraviolet scintillation light as well as free ionization electrons. The instantaneous light signal, called the S1 signal, is detected by two arrays of photomultiplier tubes (PMTs) at the top and bottom of the approximately cylindrical TPC. An electric field ( average electric drift field during SR0) is applied to the target volume to drift the ionization electrons toward the liquid xenon surface. Here, they are extracted and accelerated into a xenon gas volume above the liquid via a stronger electric field ( electric extraction field in liquid during SR0). The kinetic energy of the accelerated electrons in the gas phase is sufficient for the emission of electroluminescent light, which is proportional to the number of extracted electrons. This second light signal is referred to as the S2 signal. The distribution of the S2 signal detected by the PMTs in the top array is used to infer the position of the initial interaction in the horizontal plane parallel to the liquid surface, which defines the radial coordinate . The vertical component follows from the drift time of electrons, which is determined as the time difference between the S1 and S2 signals. This three-dimensional position reconstruction of the interaction vertex enables the discrimination of multi-site events as well as the selection of an inner fiducial volume (FV) within the TPC. This volume benefits from a particularly low background level thanks to the excellent self-shielding properties of LXe, given its high density. Moreover, the relative magnitudes of the S1 and S2 signals can be used to determine if a particle interacted with the xenon nucleus (expected from WIMPs) or xenon shell electrons (typical of the dominant sources of background), i.e. nuclear recoil (NR) or electronic recoil (ER) events. The electric drift and extraction fields are established by three parallel-wire electrodes (cathode, gate, and anode from bottom to top). To reduce wire sagging, two and four horizontal perpendicular wires support the gate and anode electrodes, respectively.
@@
The TPC is nested in an active neutron veto, which in turn is housed in an active muon veto. Both veto systems are water Cherenkov detectors. More details on the TPC, the veto systems, and other subsystems of the detector are provided in [6, 7, 8].
@@ II Modeling the detector response
Modeling the detector response to energy depositions allows for a powerful discrimination between NR signal and ER background in S1–S2 space. NR events are produced by particles scattering elastically off xenon nuclei, while in ER events, particles scatter off xenon shell electrons. The different energy loss processes involved after the two types of recoil cause ER and NR events to form separate populations in S1–S2 space, which is exploited in the WIMP DM search. The modeling is performed in the region of interest (ROI) designated for the SR0 WIMP search, defined by in [0, 100] PE (PE = photoelectron) and in [, ] PE. This ROI is selected to contain the majority of our expected signal (c.f. Figs. 4 and 5). We build the models using calibration data before unblinding the WIMP search dataset.
In this section, we describe the TPC response model to energy depositions in LXe via ER or NR interactions, which is obtained from fits to calibration data. This detector response model is separated into two parts. The first one is the empirically parametrized LXe emission model, which describes the production processes of the detectable quanta, i.e. the conversion of deposited energy into scintillation photons and ionization electrons. The second part is the detector reconstruction model, which covers the conversion from the produced photons and electrons into the observed and spacetime-dependence corrected S1 and S2 signals (cS1 and cS2). Due to its complexity, it is unfeasible in the fit to simulate processes of photon and electron propagation on a per-quantum basis. Instead, toy-Monte-Carlo (toy-MC) simulations of the detector reconstruction model are used which sample from effective maps provided either by data-driven analyses described in [11] or by simulation-driven analyses using the XENONnT Monte Carlo (MC) framework [12, 13]. All model parameters in the simulations are fit to both ER and NR calibration data. The two parts of the detector response model together with the toy-MC simulation workflow are described in Secs. II.1 and II.2, and the fit to ER and NR calibration data is detailed in Sec. II.3.
@@ II.1 Liquid Xenon Emission Model
The production of quanta from energy depositions in a xenon target is complex and lacks a comprehensive description derived from first principles. Thus, a semi-empirical parametrization of the emission model is commonly used. The parametrization used in XENONnT SR0 is similar to XENON1T [10], which is based on the Noble Element Simulation Technique (NEST) models described in [14, 15]. The simulation workflow of the emission model is described in detail in Appendices A and B, and is summarized in the following.
@@
In the SR0 WIMP search ROI, ER interactions have a much smaller ratio of produced excitons to electron-ion pairs than NR interactions (by a factor of up to 10 in the ROI), while the fraction of electron-ion pairs that recombine is similar for both interaction types given the same number of detectable quanta. The difference in exciton-to-ion ratio dominates, leading to larger ratios of S2 to S1 areas for ER interactions, which is the principle behind ER-NR discrimination in LXe TPCs. We parametrize the emission model for ER interactions as in [10], with a constant exciton-to-ion ratio, and a mean recombination fraction following the Thomas-Imel box model [16] with an additional Fermi-Dirac term. For NR interactions, the emission model follows the parametrization used in the NEST v1 model detailed in [15]. While there are newer versions of fits available in the NEST framework that use a different parametrization for NR interactions (which we refer to as NEST v2 in the following), we choose to stay with the previous version (referred to as NEST v1). The model is simpler, fits our data well, and the best-fit values and uncertainties of parameters of NEST v2 were not published at the time of the analysis.
@@ II.2 Detector Reconstruction Model
For S1 signals, the spatial dependence of geometric effects during photon propagation, the photon detection efficiency of the PMTs, and the effect of double photoelectron emission (DPE) from the PMT photocathode [17, 18] are all modeled in the simulations. For S2 signals, we model the loss of electrons along their drift path due to attachment to electronegative impurities as well as due to electric field effects close to the TPC wall, the efficiency of electron extraction from the liquid into the gas phase, and the single-electron gain of extracted electrons to detected PE via the electroluminescence process in gas. In addition to these physical processes, we also account for the influence of software reconstruction effects on the signals. For both signal types, the software reconstruction process can introduce biases and fluctuations in the recorded signal size. We also account for the S1 signal reconstruction efficiency, which vanishes at about 3 PE, as we require that at least 3 PMTs detect a photon within 50 ns around the maximal amplitude. Finally, we model the uncertainty of the event position reconstruction and the acceptances of data quality selections [11].
@@
The NR model is calibrated using an external 241AmBe neutron source (referred to as AmBe source in the following). It emits fast neutrons via the 9BeC capture reaction. Neutrons often scatter multiple times in the LXe target in so-called multi-scatter (MS) events, creating distinct energy depositions at each interaction site, thus the topologies of neutron events must be considered in the modeling. Since the neutrons travel at about 10 % of the speed of light and the S1 width is , the S1 signals from the separate energy depositions in a MS event get reconstructed into one merged S1 signal. The S2 signals might be distinguished based on the separation in spacetime between the energy depositions of the individual scatters, but can get (partially) merged if the interaction sites are close together. Due to the low drift field of in XENONnT SR0, the separation resolution between two S2 signals is reduced with respect to the design because of the increased drift time and consequently larger S2 spread due to the diffusion of electrons. This effect is accounted for in the fit of the NR model to the neutron calibration data: we first produce the photons and electrons for each energy deposition in a MS event separately, based on the emission model. While the primary scintillation photons are summed to produce the merged S1 signal, we only sum S2 signals that will become part of the largest S2 signal after going through the remaining detector reconstruction processes. This information is provided by PMT waveform simulations of neutron scatters from the AmBe source, using the Monte Carlo (MC) framework of XENONnT [12, 13] (see also Sec. III.3 for details on neutron simulations). The results of these waveform simulations are used as inputs to the toy-MC simulations and fitting framework. Data selections that remove MS events (both resolved and unresolved) are directly applied to these inputs, because their acceptances can only be determined from full simulations, and are correspondingly dropped from the selection acceptance curves for the NR fit.
@@ II.3 Fit to Calibration Data
The parameters in the full detector response model are fitted to 220Rn, 37Ar, and AmBe calibration data. The ROI selection is the same as the SR0 WIMP search region defined in Sec. II.1. All data quality selections are applied to the calibration data within the ROI. For the AmBe neutron data, an additional selection is applied using coincident gammas detected in the neutron veto. In 9BeC capture reactions in the AmBe source, 12C is left in the first excited state after neutron emission in about 50 % of the cases, estimated with neutron veto data [21]. The 12C excited state promptly decays via the emission of a 4.44 MeV gamma ray, coincident with the neutron emission. The requirement of the time coincidence within between the 4.44 MeV gamma ray detected in the neutron veto and the event in the TPC, leads to a clean calibration dataset, with only about one expected event from accidental coincidences between the two detectors in the dataset (0.05 % of the resulting sample). The events recorded in the ROI that pass (fail) the neutron veto coincidence selection are shown in red (gray) before applying any additional data quality selections in Fig. 1.
@@
In the ER emission model, the parametrization is the same as in XENON1T [10] with very wide flat priors, referred to as free priors. For the NR emission model in turn, the parameter priors are taken from [15]. A Gaussian-like distribution with different widths on either side is used as the prior of parameters with asymmetric uncertainties. However, due to the low drift field of 23 V/cm, where no literature data on NR yields exist, the validity of the field dependence in the model is unverified. Therefore, the field dependence in the emission model from [15] was modified. Two scaling parameters were freed ( and in Eq. 24 and Eq. 27 in the appendix), and two field dependence parameters ( and ) were fixed to the reported best-fit values. The widths of all other parameter priors were doubled in order to allow more freedom in the fit. Because the dependence of NR signals on the drift field has been shown to be small (see also literature measurements shown in Fig. 3), we consider this a reasonable choice.
The fit results of all emission model parameters are summarized in Tab. 2 in the appendix. No significant tension between the posteriors and the priors of any parameter is found. Suitable goodness-of-fit (GOF) tests are performed to assess whether the best-fit models adequately describe the 220Rn and AmBe calibration data. Specifically, we employ binned Poisson likelihood tests in the – space. We adopt an equiprobable binning scheme using the GOFevaluation package [24], ensuring that the number of expected events under the best-fit model is the same in each bin, . In Fig. 2 the results of the 2D – Poisson likelihood test are shown for both the ER and the NR fit, overlaid with the 220Rn and the AmBe calibration data, respectively. The resulting p-values of the tests indicate no significant discrepancy between the best-fit models and the calibration data. The few outlier events below the NR band in the right panel are modeled by accounting for S2 charge losses near the TPC walls (see also Appendix C in the appendix).
@@ III.1 WIMP Signal Model
The axial-vector structure factor of xenon is taken from [49]. Note that the uncertainty on the structure factor dominates especially for the “neutron-only” case. In the XENONnT SR0 analysis, the mean of the structure factor is used.
@@ III.2 Electronic Recoil Background Model
ER events are one of the dominant background sources in the XENONnT SR0 WIMP search, primarily originating from -decays of intrinsic radioactive contaminants such as 214Pb (a product of the 222Rn decay chain) and 85Kr. Contributions from 136Xe double- decays, solar neutrino interactions, and gamma events from the materials are also taken into account. In the ROI for the WIMP search, the total ER energy spectrum is approximately flat. The distribution of ER background events in – space is generated from the ER model fitted to 220Rn and 37Ar calibration data, as discussed in Sec. II.3. The total ER rate is fitted in the final WIMP likelihood, without applying an additional ancillary constraint in order to avoid potential systematic biases, which will be discussed in Sec. IV.
@@
along each principal component is used to quantify its importance. Here, and are the probability of a 50 GeV WIMP and ER background in the bin of cS1–cS2 space, respectively. A larger variance of means that the uncertainty of that component can be more impactful for the WIMP sensitivity. In the end, the two most important principal components are kept as nuisance parameters. These two shape parameters are constrained within the 220Rn ancillary term of the WIMP search likelihood function, as shown in Fig. 6. While adding an additional 37Ar ancillary term could further constrain the shape uncertainty at low energies, it was not included to maintain a conservative approach, as 220Rn already provides a sufficiently tight constraint on the shape uncertainty.